If the sun is 150 million km away from the earth, how long does it take sunlight to reach the earth
1
Expert's answer
2018-11-06T15:38:09-0500

We can find the time that the sunlight takes to reach the Earth from the formula:

t=d/c=(1.5∙〖10〗^11 m)/(3∙〖10〗^8 m/s)=500s.